Find measurement arc AF=(26x+15)°,find measurement arc EA


1
Expert's answer
2021-02-10T02:07:59-0500

Here arc AF=26x+15


So, as the angle is equal to the angle subtended by a diameter at center hence,


"26x+15=90\\
\Rightarrow26x=75\\
\Rightarrow x=(\frac{75}{26})^{\circ}"


Now, arc EA is:-


"(180-\frac{75}{26})^{\circ}=177.11^{\circ}"